OpenStudy (anonymous):

Is the relationship of the data below linear? Hours Worked Amount Charged 1 25 2 40 3 60 4 80

OpenStudy (anonymous):

I know all about functions and such, but I'm still fuzzy on the whole "linear or not" thing. Can someone please explain how to tell whether or not a relationship is linear? Don't just say yes or no, please.

hartnn (hartnn):

a linear function is of the form y=mx+c take any 2 points, say 3 60 2 40 can you find the equation of line using these ?(find m,c) if yes, then plug in other 2 points to check whether they lie on that line. if they do, function is linear, if they don't, the function isn't linear.

OpenStudy (anonymous):

Yeah... Didn't get that at all. Oh well.

hartnn (hartnn):

lol. ok, lets look at it another way, observe that increase in hours is constant = 1 (1+1=2.2+1=3,...) so, increase in amount charged must also be constant for a linear function. is it constant in this case ?

hartnn (hartnn):

25+ ...= 40 40 +... = 60 60+ ...= 80 if all those ... are equal, then the function is linear.